OpenStudy (anonymous):

how would you find the value of cos (-15)

OpenStudy (anonymous):

Go to calculator, and press cos then enter -15) XD

OpenStudy (anonymous):

-0.75968791285

OpenStudy (anonymous):

.9659258269

OpenStudy (anonymous):

@ColourThesis use sum and difference identities not calculator

OpenStudy (aravindg):

As cosine is an even function cos (-15)=cos (15) Now cos(15) can be written as cos (30/2)=\(\sqrt{\dfrac{1+\cos 30}{2}}\)

OpenStudy (aravindg):

I hope you can do it from there.

OpenStudy (anonymous):

@AravindG but my prof wanted me to find cos(-15) using sum and difference so how am i going to do that?

OpenStudy (anonymous):

ur Prof is mean lol

OpenStudy (anonymous):

here: http://home.windstream.net/okrebs/page101.html my prof recommended this website when we were learning this